A fifth of the larger of two consecutive interger exceeds a seventh of the smaller by 3. Find the integers?
Answer by checkley77(12844)

(x+1)/5=x/7+3
(x+1)/5-x/7=3
[7(x+1)-5x]/5*7=3
[7x+7-5x]/35=3
2x+7=3*35
2x+7=105
2x=105-7
2x=98
x=98/2
x=49 ans. for the smaller integer.
49+1=50 ans. for the larger integer.
Proof:
50/5=49/7+3
10=7+3
10=10